COULD SOMEONE SHOW ME THE STEPS TO SOLVE THE FOLLOWING PROBLEMS:
1.WHAT EXPRESSION REPRESENTS THE PRODUCT OF 3 LESS THAN TWICE X AND 2 MORE THAN THE QUANTITY 3 TIMES X?
2. IF X=-1 AND Y=2, WHAT IS THE VALUE OF THE EXPRESSION 2X^3 -3xy?
THANK YOU!!

Answer by scott8148(6628)

product means multiplication ... 3 less than twice x is 2x-3 ... 2 more than the quantity 3 times x is 3x+2

so (2x-3)(3x+2)=6x^2+4x-9x-6=6x^2-5x-6

to evaluate the expression, substitute the values of the variables .... 2(-1)^3-3(-1)(2)=-2-(-6)=4